# Understanding Pre-training and Fine-tuning from Loss Landscape Perspective

### v1

@@ 3.3 Worst-case Loss Landscape
$\bm{\delta}=\arg\max_{\bm{\delta}}L(\bm{\theta}+b\bm{\delta})\text{ s.t. }\|\bm{\delta}\|_{2}^{2}=\mathbb{E}[\|\mathcal{N}(\bm{0},\bm{I})\|_{2}^{2}]$ (3)
The constraint $\|\bm{\delta}\|_{2}^{2}=\mathbb{E}[\|\mathcal{N}(\bm{0},\bm{I})\|_{2}^{2}]$ ensures that the perturbation norm matches that used in Sec. 3.2, enabling direct comparison between the loss landscapes in Fig. 1 and Fig. 2 for each model. The hyperparameter $b$ are only designed to facilitates the identification of the worst-case $\bm{\delta}$, and does not affect the final direction. Equation (3) is solved by optimizing $L(\bm{\theta}+b\bm{\delta})$ using SGD and projecting the norm of $\bm{\delta}$ to unity at each step [47].
General Geometry. As illustrated in Fig. 2, the worst-case loss landscape resembles a cliff, regardless of the model or capability evaluated. This indicates that moving a short distance along the worst-case direction rapidly degrades all model capabilities. This phenomenon aligns with prior explanations for adversarial examples: in high-dimensional spaces, there exists a direction that causes rapid degradation, despite most directions being safe [65, 27, 47]. The parameter dimensions of large language models are significantly larger than those of earlier smaller models, making the worst-case direction potentially far more detrimental. This explains why prior adversarial fine-tuning, using only 10 samples and one epoch, can severely compromise the safety capabilities of a model [57].
@@ 3.4 SFT-case Loss Landscape
Section 3.2 demonstrates that most directions do not lead to performance degradation within a certain range, whereas Section 3.3 shows that a worst-case direction exists that rapidly degrades all capabilities. The direction of supervised fine-tuning (SFT) naturally lies between these extremes: it may not preserve all capabilities as effectively as the most-case direction, but it does not degrade as quickly as the worst-case direction.
Settings. To visualize the loss landscape along the SFT direction, we select $\bm{\delta}=\frac{\bm{\theta}_{sft}-\bm{\theta}_{0}}{\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}}\cdot\sqrt{\mathbb{E}[\|\mathcal{N}(\bm{0},\bm{I})\|_{2}^{2}]}$. This normalization and rescaling ensure that the perturbation norm matches those used in Sections 3.2 and 3.3, enabling direct comparison of the loss landscapes.
SFT Configurations. We investigate three types of supervised fine-tuning (SFT) using Qwen2.5-7B [71]. Benign fine-tuning employs a dataset similar to the original training data. We achieve this by selecting $\bm{\theta}_{0}$ as Qwen2.5-7B and $\bm{\theta}_{sft}$ as its officially fine-tuned version, Qwen2.5-7B-1M [71]. Normal fine-tuning uses a dataset with a distributional gap from the original data. We achieve this by following the setup in Section 4.4 of [58], i.e., fine-tuning on the Alpaca dataset [76] for one epoch. Adversarial fine-tuning utilizes the adversarial AdvBench dataset, fine-tuning for only 10 steps, following the setup in Section 4.2 of [58].
Results. As shown in Fig. 3(a), the loss landscape along the benign fine-tuning direction resembles the most-case landscape. It preserves safety within the most-case basin and loses capability when moving outside this basin. In Fig. 3(b), when there is a distributional gap between the SFT dataset and the original dataset, the loss landscape becomes narrower and sharper, indicating that the fine-tuning direction does not align with the most-case directions. In Fig. 3(c), when fine-tuning on the adversarial dataset, the model rapidly loses all capabilities, responding only with phrases like “Sure, here is” without providing factual answers to questions. Thus, while some SFT configurations align closely with the most-case direction (e.g., Fig. 3(a)), others deviate and degrade capabilities more rapidly (e.g., Fig. 3(b)). This variation clearly depends on the dataset and hyperparameters.
In the following section, we propose that, regardless of the dataset or the model’s sensitivity to hyperparameters, the size of the most-case basin can consistently bound performance degradation during any fine-tuning or jailbreaking attacks.
4 Theoretical Benefits
To begin with, we first formally define the size of most-case basin.
@@ 4 Theoretical Benefits
A model $f_{\bm{\theta}}$ is said to have $\sigma$-basin on benchmark $J_{\mathcal{D}}$, if it noised version $f_{\bm{\theta}+\bm{\epsilon}}$ where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ perform nearly the same as original version $f_{\bm{\theta}}$, i.e.,
$J_{\mathcal{D}}(f_{\bm{\theta}})-\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}+\bm{\epsilon}})]\leq\epsilon.$ (4)
Definition 4 is a necessary condition for a model to have most-case landscape looks like Fig. 1. In other words, if a model $f_{\bm{\theta}}$ have loss landscape that within range $\sigma$ the model perform nearly the same, then it must have $\sigma$-basin, satisfy Definition 4.
@@ 4.1 Any Alignment Can Be Bounded by Average-case Alignment
This is achieved through the concept of randomized smoothing [18, 60, 38, 14]. Since the model performs nearly identically within a $\sigma$-basin, for any input $\bm{x}$, instead of returning $f_{\bm{\theta}}(\bm{x})$, we can sample $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ and return $f_{\bm{\theta}+\bm{\epsilon}}(\bm{x})$. Thus, the benchmark value for this model is $J_{\mathcal{D}}(f_{\bm{\theta}+\bm{\epsilon}})$.
The following theorem demonstrates that, for any bounded benchmark $J_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$ [^3] Without loss of generality, any benchmark with a bounded output range can be normalized to this interval to obtain a corresponding certified bound., regardless of how sensitive $f$ is to its parameters, the smoothed model $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}+\bm{\epsilon}})]$ is at most $\frac{1}{\sqrt{2\pi}\sigma}$-Lipschitz. Consequently, when $\bm{\theta}_{0}$ is updated to $\bm{\theta}_{sft}$, the benchmark value changes by at most $\frac{1}{\sqrt{2\pi}\sigma}\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}$.
@@
(Weak Law of Randomized Smoothing [60, 13]) For any benchmark $J_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$, the function $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}+\bm{\epsilon}})]$ is at most $\frac{1}{\sqrt{2\pi}\sigma}$-Lipschitz, i.e.:
$\left|\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}_{sft}+\bm{\epsilon}})]-\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}_{0}+\bm{\epsilon}})]\right|\leq\frac{1}{\sqrt{2\pi}\sigma}\cdot\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}.$ (5)
Thus, we can bound the performance degradation as:
$\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}_{sft}+\bm{\epsilon}})]\geq\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}_{0}+\bm{\epsilon}})]-\frac{1}{\sqrt{2\pi}\sigma}\cdot\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}.$ (6)
@@
The works in [18, 60, 14] provide a stronger version of Eq. 6 by considering the maximum Lipschitz constant at each point rather than across the entire input space, as presented in Eq. 7. Consequently, Eq. 7 consistently provides a tighter bound than Eq. 6.
@@
(Strong Law of Randomized Smoothing [18, 60, 14]) For any benchmark $J_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$, we have:
$\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}_{sft}+\bm{\epsilon}})]\geq\Phi\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}_{0}+\bm{\epsilon}})]\right)-\frac{\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}}{\sigma}\right),$ (7)
@@
[…]
The term $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}+\bm{\epsilon}})]$ represents the expected benchmark value when sampling $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ and evaluating the benchmark on $f_{\bm{\theta}+\bm{\epsilon}}$. The bound in Eq. 7 provides a robust guarantee of performance stability during fine-tuning. In practice, one typically samples a single $\bm{\epsilon}$ and evaluates $f_{\bm{\theta}+\bm{\epsilon}}$, rather than computing the expectation via extensive Monte Carlo sampling. Empirically, sampling a single instance yields results comparable to multiple samples. Theoretically, the variance introduced by single-sample evaluation can be bounded using the well-known concentration phenomenon, which states that a random variable is unlikely to deviate significantly from its expectation.
@@
(Concentration of Gaussian, adapted from [66]) With probability at least $1-\delta$, we have:
$J_{\mathcal{D}}(f_{\bm{\theta}+\bm{\epsilon}})\geq\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}+\bm{\epsilon}})]-L\sigma\sqrt{2\log\frac{1}{\delta}},$ (8)
where $L$ is the Lipschitz constant of $J_{\mathcal{D}}$ with respect to $\bm{\theta}$.
By combining Eq. 7 and Theorem 4.4, we can lower bound the performance degradation when fine-tuning from $\bm{\theta}_{0}$ to $\bm{\theta}_{sft}+\bm{\epsilon}$ as follows:
Theorem 4.5
For any benchmark $J_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$, with probability at least $1-\delta$, we have:
$J_{\mathcal{D}}(f_{\bm{\theta}_{sft}+\bm{\epsilon}})\geq\Phi\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}_{0}+\bm{\epsilon}})]\right)-\frac{\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}}{\sigma}\right)-L\sigma\sqrt{2\log\frac{1}{\delta}}.$ (9)
[…]
Increasing the Basin Size Improves the Guarantee. As illustrated in Fig. 4(b), enlarging the basin of a model while preserving the performance on the original model, i.e., $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}+\bm{\epsilon}})]$, linearly strengthens the theoretical guarantee. In other words, increasing $\sigma$ results in a linear improvement in the performance degradation guarantee during fine-tuning.
@@
Figure 4: Lower bound guarantees. (a) The lower bound on the benchmark value of the smoothed fine-tuned model $J_{\mathcal{D}}(f_{\bm{\theta}_{sft}+\bm{\epsilon}})$ for varying benchmark values on the smoothed original model $\bm{\theta}_{0}$, i.e., $p_{A}:=\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}_{0}+\bm{\epsilon}})]$, with $\sigma=0.003$. (b) The lower bound on the benchmark value of the smoothed fine-tuned model $J_{\mathcal{D}}(f_{\bm{\theta}_{sft}+\bm{\epsilon}})$ for varying basin sizes $\sigma$, with $p_{A}=0.9$. (c) Histogram of L2 distances between token embeddings.
4.2 Input-space Robustness Can Be Bounded by Average-case Alignment
In this section, we propose that the size of a basin also bounds performance degradation against jailbreaking attacks [77, 70].
Intuition. Let $\bm{W}$ denote the embedding layers. Given that the embedding layers of current large language models are onto transformations (i.e., $\bm{W}$ is column full-rank) [11], the activation after a weight perturbation, $\bm{W}\bm{x}+\bm{\delta_{W}}\bm{x}$, and the activation after an input perturbation, $\bm{W}\bm{x}+\bm{W}\bm{\delta_{x}}$, can yield the same vector [74]. Thus, if the model is robust to any weight perturbation $\|\bm{\delta_{W}}\|_{2}\leq\epsilon_{\bm{W}}$, it is also robust to any input perturbation $\bm{\delta_{x}}$ such that $\bm{W}\bm{\delta_{x}}\in\{\bm{\delta_{W}}\bm{x}\mid\|\bm{\delta_{W}}\|_{2}\leq\epsilon_{\bm{W}}\}$.

### v2

@@ 3.3 Worst-case Loss Landscape
$\bm{\delta}=\arg\max_{\bm{\delta}}L(\bm{\theta}+\alpha\bm{\delta}),\;\text{ s.t. }\|\bm{\delta}\|_{2}^{2}=\mathbb{E}[\|\mathcal{N}(\bm{0},\bm{I})\|_{2}^{2}].$ (2)
The constraint $\|\bm{\delta}\|_{2}^{2}=\mathbb{E}[\|\mathcal{N}(\bm{0},\bm{I})\|_{2}^{2}]$ ensures that the perturbation norm matches that used in Figs.˜1 and 2, enabling more direct comparison for each model. Eq.˜2 is solved by optimizing $L(\bm{\theta}+b\bm{\delta})$ using SGD and projecting the norm of $\bm{\delta}$ to unity at each step (Madry et al., 2018).
General Geometry. As illustrated in Fig.˜2, the worst-case loss landscape resembles a cliff, regardless of the model or capability evaluated. This indicates that moving a short distance along the worst-case direction rapidly degrades all model capabilities. This phenomenon aligns with prior explanations for adversarial examples: in high-dimensional spaces, with high probability there exists a worst-case direction that causes rapid degradation, despite most directions being safe (Daniely & Shacham, 2020; Bubeck et al., 2021). The parameter dimensions of large language models are significantly larger than those of earlier smaller models, making the worst-case direction potentially far more detrimental. This explains why prior adversarial fine-tuning, using only 10 samples and one epoch, can severely compromise the safety capabilities of a model (Qi et al., 2024).
@@ 3.4 SFT-case Loss Landscape
Sec.˜3.2 demonstrates that most directions do not lead to performance degradation within a certain range, whereas Sec.˜3.3 shows that a worst-case direction exists that rapidly degrades all capabilities. The direction of supervised fine-tuning (SFT) naturally lies between these extremes: it may not preserve all capabilities as effectively as the most-case direction, but it does not degrade as quickly as the worst-case direction.
Settings. To visualize the loss landscape along the SFT direction, we select $\bm{\delta}=\frac{\bm{\theta}_{sft}-\bm{\theta}_{0}}{\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}}\cdot\sqrt{\mathbb{E}[\|\mathcal{N}(\bm{0},\bm{I})\|_{2}^{2}]}$. This normalization and rescaling ensure that the perturbation norm matches those used in Secs.˜3.2 and 3.3, enabling direct comparison of the loss landscapes.
SFT Configurations. We investigate three types of supervised fine-tuning using Qwen2.5-7B (Yang et al., 2024). Benign fine-tuning employs a dataset similar to the original training data. We achieve this by selecting $\bm{\theta}_{0}$ as Qwen2.5-7B and $\bm{\theta}_{sft}$ as its officially fine-tuned version, Qwen2.5-7B-1M (Yang et al., 2024). Normal fine-tuning uses a dataset with a distributional gap from the original data. We achieve this by following the setup in Section 4.4 of Qi et al. (2023), i.e., fine-tuning on the Alpaca dataset (Zheng et al., 2024) for one epoch. Adversarial fine-tuning utilizes the adversarial AdvBench dataset, fine-tuning for only 10 steps, following the setup in Qi et al. (2023).
Results. As shown in Fig.˜3(a), the loss landscape along the benign fine-tuning direction resembles the most-case landscape. It preserves safety within the most-case basin and loses capability when moving outside this basin. In Fig.˜3(b), when there is a distributional gap between the SFT dataset and the original dataset, the loss landscape becomes narrower and sharper, indicating that the fine-tuning direction does not align with the most-case directions. In Fig.˜3(c), when fine-tuning on the adversarial dataset, the model rapidly loses all capabilities, responding only with phrases like “Sure, here is” without providing factual answers to questions. Thus, while some SFT configurations align closely with the most-case direction (e.g., Fig.˜3(a)), others deviate and degrade capabilities more rapidly (e.g., Fig.˜3(b)). This variation clearly depends on the dataset and hyperparameters.
In the next section, we show that the size of the most-case basin provides a consistent bound on performance degradation, regardless of the dataset or the model’s hyperparameter sensitivity, and applies to both fine-tuning and jailbreaking attacks.
4 Theoretical Benefits of Basins
To begin, we first formally define the size of the most-case basin. We adopt a soft definition of basins, similar to the flatness definition in Andriushchenko et al. (2023):
@@ 4 Theoretical Benefits of Basins
A model $f_{\bm{\theta}}$ is said to have a $\sigma$-basin on benchmark $J_{\mathcal{D}}$, if its noised version $f_{\bm{\theta}+\bm{\epsilon}}$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$, performs nearly the same as the original version $f_{\bm{\theta}}$, i.e.,
$J_{f,\mathcal{D}}(\bm{\theta})-\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\leq\epsilon.$ (3)
Definition 3 is a necessary condition for a model to have a most-case landscape resembling that in Fig.˜1: when $\epsilon\to 0$, it becomes the strict definition of basins, as defined by $\mathbb{E}_{\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathbb{I}\{\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})=0\}]$ in Sec.˜3.2 (assuming performance does not increase under noise perturbation). Since our theoretical analysis holds for all $\epsilon$ rather than only $\epsilon=0$, we adopt this definition in this section to provide a theoretical analysis for a more general case.
@@ 4.1 Any Alignment Can Be Bounded by Average-case Alignment
This is achieved through the concept of randomized smoothing (Cohen et al., 2019; Salman et al., 2019; Lee et al., 2019). Since the model performs nearly identically within a $\sigma$-basin, for any input $\bm{x}$, instead of returning $f_{\bm{\theta}}(\bm{x})$, we can sample $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ and return $f_{\bm{\theta}+\bm{\epsilon}}(\bm{x})$. Thus, the benchmark value for this model is $J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})$. The following theorem demonstrates that, for any bounded benchmark $J_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$ [^2] Without loss of generality, any benchmark with a bounded output range can be normalized to this interval to obtain a corresponding certified bound., regardless of how sensitive $f$ is to its parameters, the smoothed model $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$ is at most $\frac{1}{\sqrt{2\pi}\sigma}$-Lipschitz. Consequently, when $\bm{\theta}_{0}$ is updated to $\bm{\theta}_{sft}$, the benchmark value changes by at most $\frac{1}{\sqrt{2\pi}\sigma}\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}$.
@@
(Weak Law of Randomized Smoothing (Salman et al., 2019)) For any benchmark $J_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$, the function $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$ is at most $\frac{1}{\sqrt{2\pi}\sigma}$-Lipschitz. Thus, we can bound the performance degradation as:
$\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]\geq\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]-\frac{1}{\sqrt{2\pi}\sigma}\cdot\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}.$ (4)
@@
Cohen et al. (2019); Salman et al. (2019) also provide a stronger version by considering the maximum Lipschitz constant at each point rather than across the entire input space, as presented in Eq.˜5. Consequently, Eq.˜5 consistently provides a tighter bound than Eq.˜4.
@@
(Strong Law of Randomized Smoothing (Cohen et al., 2019; Salman et al., 2019; Lee et al., 2019; Chen et al., 2025a)) For any benchmark $J_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$, we have:
$\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]\geq\Phi\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]\right)-\frac{\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}}{\sigma}\right),$ (5)
@@
$\Phi(x)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{x}\exp\left(-\frac{s^{2}}{2}\right)ds.$
The term $\mathbb{E}_{\bm{\epsilon}}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$ represents the expected benchmark value when sampling $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ and evaluating the benchmark on $f_{\bm{\theta}+\bm{\epsilon}}$. Eq.˜5 provides a guarantee of the expected benchmark value during fine-tuning. In practice, one typically samples a single $\bm{\epsilon}$ and evaluates $f_{\bm{\theta}+\bm{\epsilon}}$, rather than computing the expectation via extensive Monte Carlo sampling. Empirically, sampling a single instance yields results comparable to multiple samples. Theoretically, the variance introduced by sampling can be bounded using the well-known concentration phenomenon, which states that a random variable is unlikely to deviate significantly from its expectation:
@@
(Concentration of Gaussian, adapted from Wainwright (2019)) With probability at least $1-\delta$, we have:
$J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})\geq\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]-L\sigma\sqrt{2\log\frac{1}{\delta}},$ (6)
where $L$ is the Lipschitz constant of $J_{f,\mathcal{D}}$ with respect to $\bm{\theta}$.
Examples. Through hypothesis testing in Sec.˜D.4, we establish that Qwen2.5-7B (Yang et al., 2024) has a $\sigma$-basin with $\sigma=0.003$ on the safety task using the AdvBench dataset, where $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\geq 0.9$. Based on this, we derive a lower bound on safety degradation as a function of $\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}$. As shown in Fig.˜4(a), for a $\sigma=0.003$ basin, a higher performance on the original parameters, i.e., $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]$, yields a stronger guarantee on the fine-tuned parameters $\bm{\theta}_{sft}$, i.e., $J_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})$.
Increasing the Basin Size Improves the Guarantee. As illustrated in Fig.˜4(b), enlarging the basin of a model while preserving the performance on the original model, i.e., $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$, linearly strengthens the theoretical guarantee. In other words, increasing the basin size $\sigma$ results in a linear improvement in the performance degradation guarantee during fine-tuning.
@@
Figure 4: Lower bound guarantees. (a) The lower bound on the benchmark value of the smoothed fine-tuned model $J_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})$ for varying benchmark values on the smoothed original model $\bm{\theta}_{0}$, i.e., $p_{A}:=\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]$, with $\sigma=0.003$. (b) The lower bound on the benchmark value of the smoothed fine-tuned model $J_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})$ for varying basin sizes $\sigma$, with $p_{A}=0.9$. (c) Histogram of L2 distances between token embeddings.
4.2 Input-space Robustness Can Be Bounded by Most-case Alignment
In this section, we propose that the size of a basin also provides a (loose) guarantee of performance degradation against jailbreaking attacks (Zou et al., 2023; Wei et al., 2023b), by considering only the very first layer of the model.
Intuition. Let $\bm{W}$ denote the embedding layers. Given that the embedding layers of current large language models are onto transformations (i.e., $\bm{W}$ is column full-rank) (Carlini et al., 2024), the activation after a weight perturbation, $\bm{W}\bm{x}+\bm{\delta_{W}}\bm{x}$, and the activation after an input perturbation, $\bm{W}\bm{x}+\bm{W}\bm{\delta_{x}}$, can yield the same vector (Zhang et al., 2024a). Thus, if the model is robust to any weight perturbation $\|\bm{\delta_{W}}\|_{2}\leq\epsilon_{\bm{W}}$, it is also robust to any input perturbation $\bm{\delta_{x}}$ such that $\bm{W}\bm{\delta_{x}}\in\{\bm{\delta_{W}}\bm{x}\mid\|\bm{\delta_{W}}\|_{2}\leq\epsilon_{\bm{W}}\}$.